# Supernova Remnants with Mirror Dark Matter and Hyperons

### v1

@@ IV Results
Furthermore, the particle abundances in the matter depend on the entropy density and lepton fraction, particularly the hyperonic species’ appearance. In the first panel, we have a higher lepton fraction and lower $s_{B}$, therefore, the temperature of the stellar matter is lower than the intermediate stages (second and third panels) during neutrino diffusion as we shall see later (in Fig. 3). Here, the $\Lambda$-hyperons remain prominent in the stellar matter as in the other snapshots, but their appearance is relatively delayed to $R\sim 0.86R_{2.1}$, toward the core, with $5\%$ DM mass fraction compared to the other snapshorts. The other hyperonic species ($\Xi$ and $\Sigma^{-}$) are less prominent and only appear deeper in the core from $R\sim 0.66R_{2.1}$ inwards, with $\Sigma^{+}$ and $\Sigma^{0}$ not appearing in significant quantities. At $T=0$ (last panel), on the other hand, we have even fewer hyperon species in the matter, only $\Lambda$-hyperon, $\Xi^{0}$, and $\Xi^{-}$ appear in significant amounts and their appearance is further delayed than in the first stage. Here, the $\Lambda$-hyperon starts appearing at $R\sim 0.66R_{2.1}$ and the $\Xi$ species at $R\sim 0.45R_{2.1}$ for $5\%$ DM mass fraction and deeper towards the core as the DM content reduces, the $\Sigma$’s do not appear at this stage. At the intermediate stages, where $s_{B}$ increases and $Y_{L,e}$ decreases (leading to an increase in temperature), all hyperonic species emerge quickly near the star’s surface. At $R\sim 0.74R_{2.1}$ all the hyperons had appeared in the matter. The most prominent hyperon, the $\Lambda$-hyperon appear at $R\sim 0.92R_{2.1}$ in the second stage and $R\sim 0.93R_{2.1}$ in the third stage. Consequently, the production of hyperonic species is thermally favored [73, 55, 58]. A detailed analysis of the particle distribution against density can be found in [73] where the EoS was originally used to study the evolution of PNSs.
@@
We present the effect of DM on the temperature fluctuations in the stellar matter in Fig. 3. Generally, the increase in the degree of freedom of the nuclear matter leads to a decrease in temperature, since we keep the entropy density fixed whilst increasing particle degrees of freedom. This accounts for the differences in the curves of the top panel (nucleon only) and the bottom panel (nucleon plus hyperons). Additionally, the presence of DM leads to an increase in the temperature of the stellar matter. This is because the gravitational potential of the NSs increases due to the presence of the DM (as shown in Fig. 5). The NSs with more DM are more compact so the stellar matter is more compressed. The compression increases the temperature of the matter due to the release of gravitational potential energy, in accordance with the virial theorem. According to this theorem, the kinetic energy (in this case, temperature) is related to the potential energy in a gravitationally bound system ($02T+U=$ 0, where $U$ is the gravitational potential energy) [83, 84]. The first stage of the PNSs evolution $(s_{B}=1;\;Y_{L,e}=0.4)$ is when the temperature is least increased by the presence of DM, suggesting that the first instants of the birth of the PNS do not present a big difference in temperature due to DM. On the other hand, during the second and third stages, as the star loses its neutrinos and the entropy increases $(s_{B}=2)$, the presence of DM causes a greater change in the temperature profiles. Thus, a few seconds after the formation of the PNS, the DM content begins to significantly influence temperature profiles by enhancing it, especially towards the core. It is interesting to note that in nucleonic PNSs, the increase in temperature from the second to the third stage of evolution is smaller when dark matter is present. In contrast, for hyperonic stars, a higher DM content results in a greater increase in temperature between these stages. This can be attributed to the compact nature of the star when hyperons are present, which becomes even more pronounced due to the presence of DM. Thus, the presence of DM alters the thermal equilibrium state of the star, potentially prolonging its cooling phase during evolution. This effect could challenge our understanding of the star’s age and its history of thermal evolution.
@@
Figure 4: Speed of sound squared profile of a $2.1M_{\odot}$ star, considering nucleonic (top) and hyperonic (bottom) hadronic matter compositions.
@@
In Fig. 4, we calculate the square of sound speed ($c_{s}^{2}=\partial P/\partial\varepsilon$) in the star’s interior as a function of radius. This enables us to investigate the effect of the presence of DM on the EoS of the DANSs using a star of $2.1M_{\odot}$ as a medium. Generally, softer EoSs lead to a decrease in $c_{s}^{2}$ relative to stiffer ones [85, 86]. Comparing, the top panel (nucleon only) and the bottom panel (nucleon plus hyperons) we observe that an increase in degree of freedom leads to a decrease in the $c_{s}^{2}$. This is consistent with softening the EoS if the degree of freedom of the stellar matter increases [69, 72]. Comparing the curves in each panel for the different proportions of DM, we observe that DM increases the $c_{s}^{2}$. This implies that the presence of DM increases the core pressure of the DANSs. DM inside an NS is expected to increase its gravitational binding energy, which can compress the star and increase its core density. Therefore, in higher-density media, the presence of the DM increases the repulsion at short distances between the particles which could account for the rise in $c_{s}^{2}$. The presence of DM also affects the equilibrium composition of the star. This is evident in Figs. 2 and 3. In Fig. 2, the presence of DM decreases the isospin asymmetry impacting the entire particle distribution in the star. In Fig. 3, the presence of DM increases the core temperature leading to changes in the thermal equilibrium state. These changes can significantly alter the EoS, leading to an increase in $c_{s}^{2}$. High $c_{s}^{2}$ is associated with the ability of the star to resist high gravitational pressures without collapse, which could impact the maximum mass, radius, and tidal deformability thresholds [87, 88].
@@ Acknowledgements.
P.T. sincerely thanks Davood Rafiei Karkevandi for his useful discussions on dark matter. This work is part of the project INCT-FNA proc. No. 464898/2014-5. A.I. acknowledges financial support from the São Paulo State Research Foundation (FAPESP), Grant No. 2023/09545-1. T. F. thanks the financial support from the Brazilian Institutions: CNPq (Grant No. 306834/2022-7), Improvement of Higher Education Personnel CAPES (Finance Code 001), and FAPESP (Grants No. 2017/05660-0 and 2019/07767-1). D.P.M. is partially supported by CNPq (Grant No. 303490/2021-7). K.D.M. is supported by FAPESP, Grant No. 2024/01623-6. O.L. and M.D. are supported by CNPq under Grants No. 307255/2023-9, and No. 308528/2021-2, respectively, and the project No. 01565/2023-8 (Universal).

### v2

@@ IV Results
Furthermore, the particle abundances in the matter depend on the entropy density and lepton fraction, particularly the hyperonic species’ appearance. In the first panel, we have a higher lepton fraction and lower $s_{B}$, therefore, the temperature of the stellar matter is lower than the intermediate stages (second and third panels) during neutrino diffusion as we shall see later (in Fig. 4). Here, the $\Lambda$-hyperons remain prominent in the stellar matter as in the other snapshots, but their appearance is relatively delayed to $R\sim 0.86R_{2.1}$, toward the core, with $5\%$ DM mass fraction compared to the other snapshots. The other hyperonic species ($\Xi$ and $\Sigma^{-}$) are less prominent and only appear deeper in the core from $R\sim 0.66R_{2.1}$ inwards, with $\Sigma^{+}$ and $\Sigma^{0}$ not appearing in significant quantities. At $T=0$ (last panel), on the other hand, we have even fewer hyperon species in the matter, only $\Lambda$-hyperon, $\Xi^{0}$, and $\Xi^{-}$ appear in significant amounts and their appearance is further delayed than in the first stage. Here, the $\Lambda$-hyperon starts appearing at $R\sim 0.66R_{2.1}$ and the $\Xi$ species at $R\sim 0.45R_{2.1}$ for $5\%$ DM mass fraction and deeper towards the core as the DM content reduces, the $\Sigma$’s do not appear at this stage. At the intermediate stages, where $s_{B}$ increases and $Y_{L,e}$ decreases (leading to an increase in temperature), all hyperonic species emerge quickly near the star’s surface. At $R\sim 0.74R_{2.1}$ all the hyperons had appeared in the matter. The most prominent hyperon, the $\Lambda$-hyperon appear at $R\sim 0.92R_{2.1}$ in the second stage and $R\sim 0.93R_{2.1}$ in the third stage. Consequently, the production of hyperonic species is thermally favored [76, 57, 60]. A detailed analysis of the particle distribution against density can be found in [76] where the EoS was originally used to study the evolution of PNSs.
It is important to mention that in the approach adopted here (and hereafter), the gravitational mass is fixed instead of the total baryon mass ($M_{b}$). The latter quantity represents the sum of dark and visible baryon masses for DANSs or only the visible baryon mass in the absence of DM and is expected to be conserved during PNS evolution [89, 58]. As a result, the obtained stellar profiles do not necessarily correspond to the same star along the evolution lines. The corresponding values of $M_{b}$ have also been recorded in Tab. 3, and the results show that $M_{b}$ decreases as $F_{D}$ increases. This behavior has important physical implications. It suggests that DM contributes with an additional gravitational pull, allowing the NS to reach a stable configuration with fewer $M_{b}$. Since the baryonic matter (DM plus OM) becomes more compressed due to this gravitational pull, the $M_{b}$ required to reach a given gravitational mass decreases compared to the case of a lower $F_{D}$ or OM-only configuration. Additionally, since DM is assumed to be self-interacting, it contributes to the degenerate pressure within the NS. However, the overall compression remains significant, keeping the $M_{b}$ lower for stars with a higher $F_{D}$ relative to those composed entirely of OM or lower $F_{D}$. Consequently, a fewer $M_{b}$ is required to maintain hydrostatic equilibrium for a fixed gravitational mass (see Ref. [90], which discusses the impact of DM on the $M_{b}$).
@@
We present the effect of DM on the temperature distributions in the stellar matter in Fig. 4. Generally, the increase in the degree of freedom of the nuclear matter leads to a decrease in temperature, since we keep the entropy density fixed whilst increasing particle degrees of freedom. This accounts for the differences in the curves of the top panel (nucleon only) and the bottom panel (nucleon plus hyperons). Additionally, the presence of DM leads to an increase in the temperature of the stellar matter. This is because the gravitational potential of the NSs increases due to the presence of the DM (as shown in Fig. 6). The NSs with more DM are more compact so the stellar matter is more compressed. The compression increases the temperature of the matter due to the release of gravitational potential energy, in accordance with the virial theorem. According to this theorem, the kinetic energy (in this case, temperature) is related to the potential energy in a gravitationally bound system ($02T+U=$ 0, where $U$ is the gravitational potential energy) [91, 92]. The first stage of the PNSs evolution $(s_{B}=1;\;Y_{L,e}=0.4)$ is when the temperature is least increased by the presence of DM, suggesting that the first instants of the birth of the PNS do not present a big difference in temperature due to DM. On the other hand, during the second and third stages, as the star loses its neutrinos and the entropy increases $(s_{B}=2)$, the presence of DM causes a greater change in the temperature profiles. Thus, a few seconds after the formation of the PNS, the DM content begins to significantly influence temperature profiles by enhancing it, especially towards the core. It is interesting to note that in nucleonic PNSs, the increase in temperature from the second to the third stage of evolution is smaller when dark matter is present. In contrast, for hyperonic stars, a higher DM content results in a greater increase in temperature between these stages. This can be attributed to the compact nature of the star when hyperons are present, which becomes even more pronounced due to the presence of DM. Thus, the presence of DM alters the star’s thermal equilibrium state, potentially prolonging its cooling phase during evolution. This effect could challenge our understanding of the star’s age and history of thermal evolution.
@@
Figure 5: Speed of sound squared profile of a $2.1M_{\odot}$ star, considering nucleonic (top) and hyperonic (bottom) hadronic matter compositions admix DM.
Using the EoS for the OM, we can calculate the speed of sound ($c_{s}^{OM}$) directly from $(\sqrt{dP_{OM}/d\varepsilon_{OM}})|_{s_{B}}$. Thus, for each pressure value $P_{OM}$, we obtain the corresponding $c_{s}^{OM}$ for the OM. We then estimate the speed of sound $c_{s}$ (for the two fluids) at any given $P(r)$ in the DANSs within the known values of $c_{s}^{OM}$ using a cubic spline interpolation method for each point in $P_{OM}(r)$. Additionally, considering that the DM mass fraction relative to OM is small (1% and 5%), this approach serves as a reasonable first approximation in the absence of a rigorous first-principles approach. Particularly, because the two-fluid TOV system changes the internal pressure balance which is accounted for in this approach. However, an alternative approach that determines an ‘effective speed of sound’ by incorporating contributions from both matter components and studying the impact of DM on an ‘effective’ $c_{s}$ tuned by a constant parameter that regulates the DM content can be found in Ref. [93].
@@
In Fig. 5, we calculate the square of sound speed ($c_{s}^{2}=\partial P/\partial\varepsilon$) in the star’s interior as a function of radius. This enables us to investigate the effect of the presence of DM on the EoS of the DANSs using a star of $2.1M_{\odot}$ as a medium. Generally, softer EoSs lead to a decrease in $c_{s}^{2}$ relative to stiffer ones [94, 95]. Comparing, the top panel (nucleon only) and the bottom panel (nucleon plus hyperons) we observe that an increase in degree of freedom leads to a decrease in the $c_{s}^{2}$. This is consistent with softening the EoS if the degree of freedom of the stellar matter increases [72, 75]. Comparing the curves in each panel for the different proportions of DM, we observe that DM increases the $c_{s}^{2}$. This implies that the presence of DM increases the core pressure of the DANSs. DM inside an NS is expected to increase its gravitational binding energy, which can compress the star and increase its core density. Therefore, in higher-density media, the presence of the DM increases the repulsion at short distances between the particles which could account for the rise in $c_{s}^{2}$. The presence of DM also affects the equilibrium composition of the star. This is evident in Figs. 3 and 4. In Fig. 3, the presence of DM decreases the isospin asymmetry impacting the entire particle distribution in the star. In Fig. 4, the presence of DM increases the core temperature leading to changes in the thermal equilibrium state. These changes can significantly alter the EoS, leading to an increase in $c_{s}^{2}$. High $c_{s}^{2}$ is associated with the ability of the star to resist high gravitational pressures without collapse, which could impact the maximum mass, radius, and tidal deformability thresholds [96, 97].
@@ Acknowledgements.
P.T. sincerely thanks Davood Rafiei Karkevandi for his useful discussions on dark matter. This work is part of the project INCT-FNA proc. No. 464898/2014-5. A.I. acknowledges financial support from the São Paulo State Research Foundation (FAPESP), Grant No. 2023/09545-1. T. F. thanks the financial support from the Brazilian Institutions: CNPq (Grant No. 306834/2022-7), Improvement of Higher Education Personnel CAPES (Finance Code 001), and FAPESP (Grants No. 2017/05660-0 and 2019/07767-1). D.P.M. is partially supported by CNPq (Grant No. 303490/2021-7). K.D.M. is supported by FAPESP, Grant No. 2024/01623-6. O.L. and M.D. are supported by CNPq under Grants No. 307255/2023-9, and No. 308528/2021-2, respectively, and the project No. 01565/2023-8 (Universal). Kau Dalfovo Marques died just two days after the acceptance of this paper, at the age of 30. Despite her neurological disease, she fought for her life and science with excellent humor. She will be missed by everyone fortunate to work with her and will certainly remain an eternal inspiration to all of us.